### v1

# Language Models as Causal Effect Generators

Abstract
We present a framework for large language model (LLM) based data generation with controllable causal structure. In particular, we define a procedure for turning any language model and any directed acyclic graph (DAG) into a sequence-driven structural causal model (SD-SCM). Broadly speaking, an SD-SCM is a causal model with user-defined structure and LLM-defined structural equations. We characterize how an SD-SCM allows sampling from observational, interventional, and counterfactual distributions according to the desired causal structure. We then leverage this procedure to propose a new type of benchmark for causal inference methods, generating individual-level counterfactual data without needing to manually specify functional relationships between variables. We create an example benchmark consisting of thousands of datasets, and test a suite of popular estimation methods on these datasets for average, conditional average, and individual treatment effect estimation, both with and without hidden confounding. Apart from generating data, the same procedure also allows us to test for the presence of a causal effect that might be encoded in an LLM. This procedure can underpin auditing LLMs for misinformation, discrimination, or otherwise undesirable behavior. We believe SD-SCMs can serve as a useful tool in any application that would benefit from sequential data with controllable causal structure.
1 Introduction
All causal inference methods have to deal with the same fundamental underlying problem: we can only observe one state of the world at a given moment in time. If we wish to reason about what caused what, we — in some way or other — will have to reason about a hypothetical world (or worlds) that we didn’t observe. This issue is often referred to more formally as The Fundamental Problem of Causal Inference [28]: a given individual can only receive one treatment at a time, so we can only ever observe one of their potential outcomes (or counterfactuals). Overcoming this fundamental problem not only makes causal inference a challenge to perform, but also a challenge to evaluate with realistic data. Our main focus — data generation with known causal structure — is thus an important tool both for the development and the evaluation of methods that deal with causality.
A well-established strategy for causal effect estimation uses randomization of a treatment to estimate average treatment effects (ATEs) [50], capturing the effect of a treatment on average across many individuals. Because ATEs are feasible to isolate with proper study design, many realistic datasets exist with ‘true’ or at least widely-accepted ATEs [38, 15, 30]. If, instead, inferences about individual treatment effects (ITEs) are desired, making additional modeling assumptions can allow us to reason about counterfactuals at an individual level; unfortunately, so long as counterfactuals remain unobservable, these assumptions remain untestable. This means that there is no easy way to collect realistic data for ITEs. [^1] It is common in literature on individual-level causal inference to draw no distinction between ITEs and conditional average treatment effects (CATEs) [55], but we emphasize these two quantities as distinct: the CATE is the conditional expectation of the ITE, which typically will not explain all ITE variation [39]. In particular, the counterfactual outcome $\tilde{y}$ and/or treatment assignment $\tilde{t}$ must always be manually generated, even if covariates $\tilde{x}$ are based on real data (see, e.g., [42]). Our method is aimed at using LLMs to provide this type of data in a flexible manner.
Figure 1: (a) A sequence-driven structural causal model (SD-SCM) uses a language model to sample data according to any desired causal relationships between variables (blue solid arrows represent an example set of possible DAGs). Any variables whose values are sampled from the language model will potentially share the language model as a common cause (red dashed arrows), unless sampled manually, e.g., uniformly. (b) However, the effects of red dashed arrows are not necessarily significant. For example, the variable weather is a strong confounder (top) or strong collider (bottom), depending on which DAG is chosen. All structural equations are determined implicitly by whatever has already been encoded in the language model (see Section 3.3 for the full example).
Contribution.
We define a procedure for turning any language model and any DAG into a sequence-driven structural causal model (SD-SCM). We characterize in Section 3 how an SD-SCM provides access to observational, interventional, and counterfactual distributions over sequential data according to the desired causal structure, where structural equations are defined implicitly via language model (see Figure 1). In Section 4, we use SD-SCMs to create an example benchmark for causal effect estimation and test a suite of popular estimation methods across ATE, CATE, and ITE estimation. Section 5 demonstrates how this same technique can underpin the auditing of language models for (un)desirable causal effects. All benchmark datasets as well as code for SD-SCM-based data generation is available on GitHub. [^2] https://github.com/lbynum/sequence-driven-scms We believe SD-SCMs can be used across many applications as a general tool for imposing causal structure on sequentially generated data.
2 Related work
Causal inference benchmarks and evaluation.
[12] lay out four categories of commonly-used methods for semi-synthetic data generation with known causal effects: (1) simulating treatment effects using real baseline outcomes [e.g., 33]; (2) using real covariates but simulating response surfaces [e.g., 58, 20, 27]; (3) performing biased sampling of randomized data [e.g., 24, 15]; and (4) constructing (proxies of) counterfactuals and interventions from real or empirical data [e.g., 42, 23]. The paradigm of fitting models to real data and then sampling synthetic data from the fit models is common in many works [52, 45]. In this area, the most closely related works to ours in spirit are those that fit generative models to real datasets such that treatments, outcomes, and covariates — in effect, entirely new datasets — can be sampled, such as [2] and [45]. While such methods are similar in that they rely on generative models, they are fundamentally different from ours, as they are based on individual datasets that already exist (and already have a fixed causal structure), rather than allowing for arbitrary causal structures to be imagined by a user and then parameterized by a generative model. Our setup is akin to a high-fidelity simulation environment [e.g., 44] that provides empirical counterfactual data [23], but without needing to manually design all aspects of the simulation, and in a manner that is instead based on natural language. This work is also loosely related to methods that parameterize structural causal models (SCMs) with generative models or other deep learning components, such as [47, 51], but such methods are geared towards counterfactual inference and learning causal relationships from existing data, rather than flexible data generation.
Language models and causal inference.
Our work is not the first to suggest that language models can generate outputs that have casual structure. Many works aim to augment language models with the ability to generate counterfactual data [7, 40, 6, 25, 22]. Counterfactuals and causal reasoning are useful across various natural language processing (NLP) tasks, making this capability of particular interest for ongoing LLM research [57], and language models with causal reasoning capabilities have a wide variety of applications both within and beyond NLP [54, 31, 59, 41, 18, 37, 32, 22]. We are also not the first to point out that counterfactual data generation with language models is useful for understanding the internal ‘world model’ constructed by an LLM and auditing for bias [7, 21]. However, our focus in this work is fundamentally different: rather than explore what language models can do with causal reasoning and causal structure across different types of tasks, we instead focus on how to generate data given a desired causal structure. This capability has important use-cases for downstream tasks like the one we demonstrate here — generating treatment effects to benchmark effect estimation methods. But more broadly, we provide a generalization of how sequence data and structural causal models can be combined in order to flexibly generate observational, interventional, and counterfactual data for whatever purpose it might be useful.
3 Controlled causal data generation via language model
We first introduce notation preliminaries in Section 3.1 before defining our procedure in Section 3.2.
3.1 Preliminaries
Let lowercase letter with tilde $\tilde{v}$ denote a random variable, where $\tilde{v}=v$ denotes the value it obtains. Let boldface capital letter $\mathbf{V}=\{\tilde{v}_{1},\ldots,\tilde{v}_{n}\}$ denote a set of variables with value $\mathbf{V}=\mathbf{v}$, capital $P_{\tilde{v}}$ denote the cumulative distribution function of $\tilde{v}$, and lowercase $p_{\tilde{v}}$ denote the density (or mass) function. Let $P_{\tilde{v}\mid\tilde{x}=x}$ denote the conditional distribution of $\tilde{v}$ given $\tilde{x}=x$ and $P_{\tilde{v}\mid\tilde{x}}$ denote the collection of $P_{\tilde{v}\mid\tilde{x}=x}$ for all $x$. A sequence, or string, is an ordered collection of tokens. We represent this either as a tuple (e.g., sequence $v=(w_{1},\ldots,w_{T})$ has tokens $w_{t}$), or interchangeably as a single string (e.g., $v=w_{1:T}\equiv\bigoplus_{t=1}^{T}w_{t}$, where $\oplus$ represents string concatenation).
Definition 1 (Language model)
Given a vocabulary $\mathbb{V}$ of possible tokens, we define a language model $\mathcal{P}$ as a joint distribution over any sequence of tokens $v=(w_{1},\ldots,w_{T})\in\bigtimes_{t=1}^{T}\mathbb{V}$, where $\mathcal{P}(v)=\prod_{t=1}^{T}\mathcal{P}(w_{t}\mid w_{1:(t-1)})$.
Definition 2 (Structural causal model)
We define a structural causal model (SCM)as a 4-tuple $\mathfrak{C}=(\mathbf{V},\mathbf{U},\mathbf{F},P_{\mathbf{U}})$. In this tuple, $\mathbf{V}$ is a set of observed variables, $\mathbf{U}$ a set of unobserved (exogenous) variables, $\mathbf{F}$ a set of functions $\{f_{i}\}_{i=1}^{|\mathbf{V}|}$ for each $\tilde{v}_{i}\in\mathbf{V}$ such that $\tilde{v}_{i}=f_{i}(\mathbf{PA}_{i},\mathbf{U}_{i})$ where $\mathbf{PA}_{i}\subseteq\mathbf{V}\setminus\{\tilde{v}_{i}\}$ represents the causal parents of $\tilde{v}_{i}$ and $\mathbf{U}_{i}\subseteq\mathbf{U}$, and $P_{\mathbf{U}}$ a distribution over $\mathbf{U}$. A causal model can be represented visually as a directed acyclic graph (DAG)with nodes for $\mathbf{U},\mathbf{V}$ and directed edges for $\mathbf{F}$. SCMs entail an observational distribution $P^{\mathfrak{C}}$ across variables $\mathbf{V}\cup\mathbf{U}$.
Definition 3 (Interventional distribution)
An SCM $\mathfrak{C}$ also entails the distribution of any subset of variables in $\mathbf{V}\cup\mathbf{U}$ following atomic intervention $I=\text{do}\left(\tilde{v}_{i}:=v\right)$, which replaces the structural mechanism $f_{i}$ with fixed value $v$. Interventions can also be extended to general modifications of $f_{i}$. We denote an SCM after intervention $I$ as $\mathfrak{C}^{\text{do}(I)}$ and the resulting distribution as $P^{\mathfrak{C};\text{do}(I)}$. [^3] Our notational conventions for interventional and counterfactual distributions follow [48].
Counterfactual distributions are computed in a similar fashion, but first conditioning $P_{\mathbf{U}}$ on a particular context before performing an intervention. Where ambiguous, we use an asterisk to denote counterfactual versions $\mathbf{V}^{*}$ of factual variables $\mathbf{V}$ [4].
Definition 4 (Counterfactual distribution)
Counterfactual variable $\mathbf{Y}^{*}$ given a factual observation $\mathbf{z}$ and intervention $\text{do}(I)$ (where $\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}$) can be computed via a three-step procedure often referred to as ‘abduction, action, prediction.’ Abduction uses observed evidence to obtain $P_{\mathbf{U}\mid\mathbf{Z}=\mathbf{z}}$ from $P_{\mathbf{U}}$. Action performs intervention $\text{do}(I)$ to obtain modified SCM $\mathfrak{C}^{\text{do}(I)}$. Prediction computes the probability of $\mathbf{Y}^{*}$ from $\mathfrak{C}^{\text{do}(I)}$ and $P_{\mathbf{U}\mid\mathbf{Z}=\mathbf{z}}$. For general intervention $I$ and observed assignment $\mathbf{Z}=\mathbf{z}$, we denote the counterfactual distribution $P^{\mathfrak{C}\mid\mathbf{Z}=\mathbf{z};\text{do}(I)}$.
3.2 Sequence-driven structural causal models (SD-SCMs)
Consider a collection of ordered random variables $(\tilde{v}_{1},\tilde{v}_{2},\tilde{v}_{3},\ldots)$, whose sample spaces $\Omega_{\tilde{v}_{i}}$ each consist of sets of sequences. We define $\tilde{v}_{1:m}\equiv\bigoplus_{t=1}^{m}\tilde{v}_{t}$ as the concatenation of the sequences themselves. The sample space for the concatenation of sequences is the cartesian product of the constituent sample spaces $\bigtimes_{t=1}^{m}\Omega_{\tilde{v}_{t}}$. For brevity, we will use the term sequence variable to refer to a random variable whose sample space is a set of sequences. Two straightforward abstractions allow us to define SD-SCMs: domain-restricted sampling and parent-only concatenation.
Definition 5 (Domain-restricted sampling)
Given language model $\mathcal{P}$, some prior inputs $C$, and a sequence variable $\tilde{v}_{i}$ with sample space $\Omega_{\tilde{v}_{i}}$, domain-restricted sampling defines a distribution $\mathcal{P}_{\tilde{v}_{i}\mid C}$ over sample space $\Omega_{\tilde{v}_{i}}$ simply by tabulating and subsequently normalizing the output probabilities for each possible $v\in\Omega_{\tilde{v}_{i}}$ conditional on prior inputs $C$: $\mathcal{P}_{\tilde{v}_{i}\mid C}(v)\equiv\frac{\mathcal{P}(v\mid C)}{\sum_{v^{\prime}\in\Omega_{\tilde{v}_{i}}}\mathcal{P}(v^{\prime}\mid C)}.$
Definition 6 (Parent-only concatenation)
Given DAG $\mathcal{G}$ over $m$ sequence variables $(\tilde{v}_{1},\ldots,\tilde{v}_{m})$ and a topological ordering $\tau$ consistent with $\mathcal{G}$, parent-only concatenation defines $\left(\tilde{v}_{i}\mid\mathbf{PA}_{i}\right)\equiv\bigoplus_{t\in\mathbf{PA}_{i}}\tilde{v}_{t}$, where $\mathbf{PA}_{i}$ are the parents of $\tilde{v}_{i}$ in $\mathcal{G}$ ordered according to $\tau$.
Given a DAG $\mathcal{G}$ and a language model $\mathcal{P}$, a corresponding sequence-driven SCM defines a sample space of sequences for each variable in $\mathcal{G}$ and provides access to observational, interventional, and counterfactual distributions as follows.
Definition 7 (Sequence-driven structural causal model (SD-SCM))
We define a sequence-driven structural causal model as a 5-tuple $\mathfrak{B}=(\mathbf{V},\mathbf{U},\mathcal{G},\mathcal{P},\tau)$, where
  • $\mathbf{V}$ is a set of finite-domain endogenous/observed sequence variables and $\mathbf{U}$ a set of finite-domain exogenous/unobserved sequence variables;
  • $\mathcal{G}$ is a DAG over the variables $\tilde{x}_{i}$ in $\mathbf{V}\cup\mathbf{U}$ where $\mathbf{PA}_{i}\subseteq\left(\mathbf{V}\cup\mathbf{U}\right)\setminus\{\tilde{x}_{i}\}$;
  • $\mathcal{P}$ is a language model trained on prior inputs $C$ whose vocabulary $\mathbb{V}$ contains all tokens used in $\Omega_{\mathbf{V}},\Omega_{\mathbf{U}}$; and
  • $\tau$ is an arbitrary fixed topological ordering of $\mathbf{V}\cup\mathbf{U}$ consistent with $\mathcal{G}$.
An SD-SCM uses $\mathcal{P}$ to define an observational distribution over the variables in $\mathbf{V}\cup\mathbf{U}$ that factorizes according to $\mathcal{G}$ via domain-restricted ancestral sampling and parent-only concatenation with $\tau$: $P^{\mathfrak{B}}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{PA}_{t}}$.
The key difference between an SCM and an SD-SCM is that all variables have at least one common ancestor — the prior inputs $C$ that were used to train the language model. As with the observational distribution, domain-restricted ancestral sampling and parent-only concatenation also allow us to define interventional and counterfactual distributions for an SD-SCM.
Definition 8 (Sequence-driven interventional distribution)
An SD-SCM $\mathfrak{B}$ entails the distribution of any subset of variables in $\mathbf{V}\cup\mathbf{U}$ following intervention $I=\text{do}\left(\tilde{v}_{i}=v\right)$ by replacing variable $\tilde{v}_{i}$ with value $v$, and otherwise sampling in the same manner. As with an SCM, we denote an SD-SCM after intervention $I$ as $\mathfrak{B}^{\text{do}(I)}$ and the resulting interventional distribution as $P^{\mathfrak{B};\text{do}(I)}$. This is computed for intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\tilde{v}_{i}=v,\mathbf{PA}^{\prime}_{t}}$, where $\mathbf{PA}^{\prime}_{t}$ denotes the parents of $\tilde{x}_{t}$, excluding $\tilde{v}_{i}$ if present.
In order to admit unique answers to counterfactual queries, we define abduction for an SD-SCM given evidence $\mathbf{Z}=\mathbf{z}$ as the setting of values $\mathbf{U}=\mathbf{u}$ and any evidence upstream of the intervention, rather than a distribution $P_{\mathbf{U}\mid\mathbf{Z}=\mathbf{z}}$. [^4] Other choices can be explored here, which we leave to future extensions of SD-SCMs. In order to obtain such values $\mathbf{u}$, one needs access to more than just the observed data and language model $\mathcal{P}$ — obtaining $\mathbf{u}$ requires performing bookkeeping during the data generation process. This is a restatement of the fact that computing point counterfactuals in SCMs requires causal mechanisms that are invertible with respect to the noise variables in order to uniquely reconstruct the noise that produced a given observation. Because our primary application of SD-SCMs in this work is data generation, such bookkeeping is possible in all our use cases.
Definition 9 (Sequence-driven counterfactual distribution)
Counterfactual sequence variable $\mathbf{Y}^{*}$ given factual evidence $\mathbf{z}$ and intervention $\text{do}(\mathbf{X}=\mathbf{x})$ (where $\mathbf{X},\mathbf{Y},\mathbf{Z}\subseteq\mathbf{V}$) can be computed for an SD-SCM $\mathfrak{B}$ whenever the exogenous setting $\mathbf{u}=\{u_{1},u_{2},\ldots,u_{|\mathbf{U}|}\}$ that generated evidence $\mathbf{z}$ is known. As with an SCM, for intervention $I$ and observed $\mathbf{Z}=\mathbf{z}$, we denote the counterfactual distribution $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(I)}$. This is computed for evidence $\mathbf{z}$, exogenous conditions $\mathbf{u}$, and intervention $\text{do}(\tilde{v}_{i}=v)$ as follows: $P^{\mathfrak{B}\mid\mathbf{Z}=\mathbf{z};\text{do}(\tilde{v}_{i}=v)}\equiv\prod_{\tilde{x}_{t}\in\tau}\mathcal{P}_{\tilde{x}_{t}\mid C,\mathbf{U}=\mathbf{u},\mathbf{Z}^{\prime}=\mathbf{z}^{\prime},\tilde{v}_{i}=v,\mathbf{PA}^{\prime\prime}_{t}}$, where $\mathbf{Z}^{\prime}\subseteq\mathbf{Z}$ contains all ancestors of $\tilde{v}_{i}$ present in $\mathbf{Z}$, and $\mathbf{PA}^{\prime\prime}_{t}=\mathbf{PA}_{t}\setminus\left(\mathbf{U}\cup\mathbf{Z}^{\prime}\cup\{\tilde{v}_{i}\}\right)$.
In summary, data can be generated from an SD-SCM by domain-restricted forward sampling variables in topological order, and, with adequate bookkeeping, both interventional and counterfactual samples can also be drawn. The key difficulty in generating data this way that is also useful for benchmarking causal inference methods is to ensure it has meaningful structure. In short, it is easy to generate data, but more difficult to generate useful data. The reason for this challenge is that we do not directly specify the structural equations; rather, the structural equations are determined by whatever the language model $\mathcal{P}$ has already encoded.
3.3 Example SD-SCM: going to the gym to improve your marathon time
Let sequence variables $\mathbf{U}=\{\tilde{u}_{1},\tilde{u_{2}}\}$ and $\mathbf{V}=\{\tilde{w}=$weather, $\tilde{g}=$gymOrRun, $\tilde{m}=$ marathon}}\}} be defined over the following sample spaces.
[…]
For this example, we use GPT-2 [49] as language model $\mathcal{P}$ and consider whether going for a run outside ($\tilde{g}=1$) will worsen marathon running times ($\tilde{m}=1$), compared to going to the gym ($\tilde{g}=0$). Figure 1 demonstrates two SD-SCMs ($\mathfrak{B}_{1}$ and $\mathfrak{B}_{2}$), corresponding to two different DAGs, $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. $\mathcal{G}_{1}$ makes weather a confounder: $\tilde{u}_{1}\rightarrow\texttt{gymOrRun}\leftarrow\texttt{weather}\rightarrow\texttt{marathon}\leftarrow\tilde{u}_{2}.$ $\mathcal{G}_{2}$ instead makes weather a collider: $\tilde{u}_{1}\rightarrow\texttt{gymOrRun}\rightarrow\texttt{weather}\leftarrow\texttt{marathon}\leftarrow\tilde{u}_{2}.$ Therefore, $P^{\mathfrak{B}_{1}}$ factorizes according to $\mathcal{G}_{1}$ as $P^{\mathfrak{B}_{1}}=\mathcal{P}_{\tilde{g}\mid C,\tilde{w},\tilde{u}_{1}}\mathcal{P}_{\tilde{m}\mid C,\tilde{w},\tilde{u}_{2}}\mathcal{P}_{\tilde{w}\mid C}\mathcal{P}_{\tilde{u}_{1}\mid C}P_{\tilde{u}_{2}},$ where we use a uniform distribution for $P_{\tilde{u}_{2}}$, as shown in Figure 1(a). $P^{\mathfrak{B}_{2}}$ instead factorizes according to $\mathcal{G}_{2}$: $P^{\mathfrak{B}_{2}}=\mathcal{P}_{\tilde{w}\mid C,\tilde{g},\tilde{m}}\mathcal{P}_{\tilde{g}\mid C,\tilde{u}_{1}}\mathcal{P}_{\tilde{m}\mid C,\tilde{u}_{2}}\mathcal{P}_{\tilde{u}_{1}\mid C}P_{\tilde{u}_{2}}$. Under $P^{\mathfrak{B}_{1}}$, we would
Figure 2: Estimation error for random forest predictions of the SATE, with and without including weather $\tilde{w}$.
expect the magnitude of correlation $\rho$ between $Pr(\tilde{g}=1)$ (the propensity of running) and $Pr(\tilde{m}=1)$ (the propensity of marathon worsening) to decrease if we condition on confounder $\tilde{w}$. Under $P^{\mathfrak{B}_{2}}$, where $\tilde{w}$ is a collider, we would instead expect the magnitude of $\rho$ to increase. Figure 1(b) shows that this is indeed the case, computed across 1000 samples drawn from each observational distribution. We can see how this would, in turn, impact causal effect estimation by drawing the corresponding 1000 counterfactual samples from $P^{\mathfrak{B}\mid\mathbf{V}=\mathbf{v};\text{do}(\tilde{g}=0)}$ and/or $P^{\mathfrak{B}\mid\mathbf{V}=\mathbf{v};\text{do}(\tilde{g}=1)}$ for each observed unit and comparing estimates of, e.g., the sample average treatment effect (SATE) to the true SATE. The SATE describes the average effect over the given sample, which we can calculate directly as the observed units’ average ITE. Letting $\tilde{y}_{i}(1),\tilde{y}_{i}(0)$ denote the outcomes of the $i$-th unit given $\text{do}(\tilde{g}=1)$ and $\text{do}(\tilde{g}=0)$, respectively, we calculate the SATE across $n$ units as $\frac{1}{n}\sum_{i=1}^{n}\left(\tilde{y}_{i}(1)-\tilde{y}_{i}(0)\right)$. Figure 2 shows prediction error (in standard deviation units) when using a random forest to predict the SATE with $P(\texttt{marathon}=1)$ as the outcome, either using treatment $\tilde{g}$ as the only covariate, or using both $\tilde{g}$ and $\tilde{w}$. As we would expect, including a confounder leads to more accurate effect estimation, while including a collider does not.
4 Generating a benchmark for causal effect estimation
Figure 3: (a) A useful DAG structure for an SD-SCM-generated estimation benchmark. (b) Visual depiction of the structure in (a) used to create the breast cancer SD-SCM in Section 4.1.
In approaching the design of an SD-SCM-generated benchmark, we focus on the fully sequential DAG structure shown in Figure LABEL:fig:benchmark_dag. Exogenous variables $\mathbf{U}$ precede covariates $\mathbf{X}$, which in turn precede treatment $\tilde{t}$. All variables precede outcome $\tilde{y}$. Recall that the presence of an edge in a DAG allows for the possibility of a relationship, but it is the structural equations that determine whether or not a given relationship is meaningful. The strongest assumptions encoded by a DAG, then, are those edges that are not present. Our goal in this work is to have a language model $\mathcal{P}$ make as many ‘decisions’ about the data generating process as possible. We thus choose this structure as a means of letting $\mathcal{P}$ define whichever structural equations are meaningful or not, and focus on the edge $\tilde{t}\rightarrow\tilde{y}$ as the target for effect estimation. The key criterion we consider for a useful effect estimation benchmark is that the datasets we generate require the use of causal reasoning (e.g., controlling for confounding) in order to recover the effect of $\tilde{t}$ on $\tilde{y}$. Specifically, we aim to generate data for which the observational and interventional distributions are different, i.e., $P^{\mathfrak{B}}_{\tilde{y}\mid\tilde{t}=t}\neq P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$. In other words, there should be confounders present within $\mathbf{U}\cup\mathbf{X}$. This criterion is not directly in our control given a fixed language model $\mathcal{P}$ [^5] We discuss applications of this same idea to training a model in Section 6. However, even with fixed $\mathcal{P}$ and a fixed DAG structure, we find we are able to achieve it through our choice of sample spaces $\Omega_{\mathbf{U}},\Omega_{\mathbf{X}},\Omega_{\tilde{t}},\Omega_{\tilde{y}}$.
4.1 Breast cancer SD-SCM
We define an SD-SCM over 14 variables in order to explore the effect of a tumor’s PD-L1 expression levels on different breast cancer therapy plans. Rather than reflect medically accurate causal relationships, our goal with this SD-SCM is instead to induce causal structure that can challenge estimation algorithms. Covariates in the breast cancer SD-SCMs are defined in full detail in Appendix A and correspond to the DAG in Figure LABEL:fig:bcancer_dag. For each covariate, 10 different phrasings are considered, resulting in a sample space of $10^{14}$ possible sequences. [^6] Language models are also used to generate the phrasings, but we leave full automation of this process to future work. For example, for the covariate $\tilde{u}_{1}$ that represents ‘age,’ with possible values $\Omega_{\tilde{u}_{1}}=(25,35,45,55,65,75,85)$, two possible phrasings are:
  1. 1.
    A $\tilde{u}_{1}$-year-old woman seeks consultation at the oncology clinic after being recently diagnosed with invasive breast cancer.
  2. 2.
    At the oncology clinic, a $\tilde{u}_{1}$-year-old woman is evaluated following a recent diagnosis of invasive breast carcinoma.
We consider 50 different SD-SCM variations, where the sample space for a given SD-SCM is defined by choosing a randomly sampled phrasing from among the possible phrasings for each of the 14 covariates. Then, for each of the 50 SD-SCMs, 20 datasets (each of size 1000) are sampled, for a total of 1000 datasets per language model. In this section, we show results for GPT-2 [49] and Llama-3-8b [17], but we emphasize that the language model is a fully modular component, and thus other language models can be used. Figure 4 shows plots of the features (LABEL:fig:bcancer_hist), propensity scores (LABEL:fig:bcancer_propensity), and ITEs (LABEL:fig:bcancer_ites_binary, LABEL:fig:bcancer_ites) for a single generated dataset using Llama-3-8b. Because the outcome $\tilde{y}$ has $|\Omega_{\tilde{y}}|=4$ possible values, we can consider several possible outcomes, including the observed outcome (categorical), log probabilities for each outcome value (continuous), or probabilities for each outcome value (continuous). This creates, in effect, nine possible targets for each dataset. For benchmarking purposes, we find using probabilities and/or log probabilities as the outcome to be the most useful — there is frequently an effect (even at the individual level) in probability space, even if the sampled outcomes do not change. Comparing Figure LABEL:fig:bcancer_ites_binary to Figure LABEL:fig:bcancer_ites demonstrates this, where for the 400 or so observations where the categorical ITE is zero, the continuous ITE is instead nonzero. Figure LABEL:fig:bcancer_ites also demonstrates that we are able to satisfy our key criterion stated above: the observational distribution $P_{\tilde{y}\mid\tilde{t}=t}^{\mathfrak{B}}$ and interventional distribution $P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$ are different enough that the SATE and the observed mean difference in outcomes between the treatment and control group are not only different in value, but they also disagree in sign. This is particularly meaningful in a causal inference setting — the treatment appears to lower the outcome, when in fact, its effect is to increase the outcome. We find that how similar $P_{\tilde{y}}^{\mathfrak{B};\text{do}(\tilde{t}=t)}$ is to $P_{\tilde{y}\mid\tilde{t}=t}^{\mathfrak{B}}$ varies across SD-SCMs, explored further by comparing the performance of observational versus casual estimation approaches.
Figure 4: An example dataset generated by the breast cancer SD-SCM using Llama-3-8b, showing (a) features, (b) propensity scores, (c) categorical outcome ITEs, and (d) continuous outcome ITEs.
4.2 Effect estimation results
We compare the performance of several treatment effect estimation algorithms on this example benchmark. As a naive baseline, ordinary least squares using only the treatment $\tilde{t}$ is considered (T-Only OLS). Against this baseline, we consider several causal inference methods of different types, including the causal forest [56, 3] (CausalForest) and two double machine learning methods for CATE estimation, one linear (LinearDML) and one non-parametric (ForestDML) [9, 3, 46, 8, 19, 43, 5]. For comparison we include also two doubly robust meta-learning methods [35], again, one linear (LinearDR) and one non-parametric (ForestDR). We add also Bayesian additive regression trees (BART) [27, 11] as a widely-used Bayesian non-parametric example. As points of reference for NN-based CATE estimation methods, we include an NN-based T-learner (TNet), and the NN-based TARNet [53]. Additional baselines include a random forest baseline (RF) that fits a single response surface and directly predicts treatment effects for each unit, and a linear regression baseline (LinReg) that takes the conditional mean difference (the fit coefficient on $\tilde{t}$) to be the effect. Finally, we include two methods that target ITEs specifically. One method uses BART posterior draws specifically for ITEs instead of CATEs (BART-ITE), and the other is conformalized counterfactual quantile regression (CQR) [39], which provides conformal inference-based interval estimates of ITEs.
All methods are fit using the default settings of their publicly-available implementations. The causal forest, DML, and DR implementations are provided by [5], the BART methods by [16], the NN-based methods by [12, 14, 13] and CQR by [39]. [^7] All code to reproduce this benchmark is available at https://github.com/lbynum/sequence-driven-scms. While additional hyperparameter tuning and model optimization could be performed for several methods on a case-by-case basis, this section demonstrates what estimation results we get off-the-shelf.
4.2.1 Average treatment effects
Table 1: SATE prediction results for methods that directly target ATEs.
Method $\mathcal{P}=$ GPT-2 $\mathcal{P}=$ Llama-3-8b
$R^{2}$ RMSE $R^{2}$ RMSE
All Cov. Hidden All Cov. Hidden All Cov. Hidden All Cov. Hidden
T-Only OLS 0.6047 0.6047 0.0172 0.0172 0.5082 0.5082 0.0091 0.0091
BART 0.9999 0.8794 0.0003 0.0095 0.9967 0.8476 0.0007 0.0051
ForestDML 0.9941 0.8686 0.0021 0.0099 0.9608 0.8129 0.0026 0.0056
ForestDR $\leq 0$ $\leq 0$ 6.4268 29.0210 0.9581 0.8179 0.0027 0.0055
LinReg 0.9146 0.8646 0.0080 0.0101 0.6538 0.5599 0.0076 0.0086
LinearDML 0.979 0.8655 0.0040 0.0100 0.9608 0.8216 0.0026 0.0055
LinearDR $\leq 0$ $\leq 0$ 11.0736 29.5414 0.9589 0.8176 0.0026 0.0055
RF 0.976 0.8766 0.0042 0.0096 0.8122 0.8295 0.0056 0.0054
For all implementations that directly support ATE estimation, we report the $R^{2}$ and root-mean-squared-error (RMSE) across the 1000 datasets for each language model in two settings, using $\text{logP}(\tilde{y}=0)$ as the outcome. The first setting is with estimation using all 14 covariates (all 12 confounders, the treatment, and the outcome). This is denoted All Cov. in Table 1. The second setting is with the variables $\mathbf{U}=\{\tilde{u}_{1},\tilde{u}_{2},\tilde{u}_{3},\tilde{u}_{4}\}=$age, medical conditions, medication, menopausal status} hidden, denoted Hidden. Table 1 shows that ATE estimation is more or less challenging depending on which language model is used. In this case, Llama-3-8b produces ATEs that are more challenging to estimate, with the exception of GPT-2 for the doubly robust methods, whose $R^{2}$ and RMSE suffer significantly due to several large outlying estimates. Across all methods, performance tends to drop significantly in the ‘Hidden’ setting, suggesting that $\mathbf{U}$ are indeed hidden confounders. We refer to this setting henceforth as ‘Hidden Confounding.’ Across methods, BART shows the strongest performance in all settings in Table 1.
As discussed in Section 1, a particular benefit of SD-SCM-based data generation is access to individual-level potential outcomes rather than ATEs only, allowing for comparisons of performance on CATE and ITE estimation.
4.2.2 Conditional and individual treatment effects
Figure 5: $R^{2}$ values (clipped at zero) across all methods that estimate CATEs.
Figure 5 shows $R^{2}$ values (clipped at zero) across all methods that provide point estimates for CATEs. When all covariates are observed, BART does significantly better explaining CATE variation, while DML and DR methods do as well at times but with a much lower average. However, CATE estimation becomes much more challenging for all methods under hidden confounding. The difference in effect estimation difficulty between Llama-3-8b and GPT-2 is also more noticeable for CATEs than it was for ATEs. Figures 9 and 10 in Appendix B show the same results in terms of PEHE (Precision in Estimating Heterogeneous Effects) [27], revealing that with no clipping, DR methods show large outliers with GPT-2, and NN methods show large outliers instead with Llama-3-8b.
Figure 6: CATE empirical coverage ($\alpha=0.05$) and interval width (in outcome standard deviation units) for methods that provide intervals. Nominal coverage of 95% is indicated by the red line.
Figure 7: Interval estimates from BART versus BART-ITE across two example datasets, (a) and (b).
When the ITE varies due to any covariates not conditioned on in the CATE, the two quantities are distinct. Thus, the Hidden Confounding setting for CATE estimation shows us how CATE estimators perform when the target is instead the ITE. In this setting, where estimation is difficult and no methods perform well, uncertainty is especially important. Figure 6 shows empirical coverage results for all estimators that provide intervals. Empirical coverage is under nominal for all methods that target CATE in the setting with all covariates. Hidden confounding generally increases uncertainty, but this does make coverage closer to nominal, especially for LinearDML and LinearDR. CQR remains at or above nominal coverage, but with much wider intervals than the other methods. Interestingly, BART for CATE achieves higher median coverage of the ITE than BART-ITE, but with a much larger tail of poor coverage. BART-ITE, by contrast, has much less variable coverage in the ITE setting, but at the cost of much wider intervals. Figure 7 demonstrates this further with intervals for BART targeting the CATE versus BART-ITE across two example datasets. When all covariates are observed (top row), there are cases where intervals from either method are informative about individual-level effects. However, under hidden confounding (bottom row), the tighter intervals of BART targeting the CATE are overconfident with variable coverage, and the wider intervals of BART-ITE are so wide as to be vacuous, even if we want to recover just the ranking of the ITEs.
Takeaways.
We summarize a few takeaways for off-the-shelf estimation performance with these example SD-SCMs. The first is that less complicated methods are often able to perform well. Second, in a real-world setting, where a method might often be used with its default parameters, stability can be important (e.g., the DR and NN method performance suffers often due to lack of stability). The third takeaway is that hidden confounding has a big impact, across all methods. Even methods that perform particularly well with all covariates (like BART) suffer significantly under hidden confounding. Finally, ITE intervals can be unstable and/or vacuous for decision making, especially under hidden confounding, and should thus be used carefully.
5 Auditing language models for (un)desirable causal effects
The same framework we use to generate causal effects and benchmark effect estimation methods can allow us to inspect what causal information has been encoded semantically in an LLM. Our collection of breast cancer SD-SCMs are already set up to explore the effect of PD-L1 on chosen therapy plans, while allowing us to marginalize out an important source of variability: phrasing. Essentially, this amounts to reverse engineering the decision-making process of clinicians, as learned
Figure 8: SATEs across all breast cancer SD-SCMs for outcome $P(\texttt{therapy}=2)$.
from whatever data the language model were trained on. Recall from Section 4.1 that there are nine possible outcomes we could explore in the breast cancer example. Figure 8 shows one example where the two language models strongly disagree on what the causal effect is. The effect in this case is the change in probability of choosing the therapy plan “start a regimen of trastuzumab and pertuzumab” (shown in standard deviation units). GPT-2 has encoded that on average, an increase in PD-L1 expression levels has neither a positive nor negative impact on whether or not this therapy will be chosen. However, Llama-3-8b has encoded instead that an increase in PD-L1 always increases the likelihood of this therapy plan. This discrepancy indicates that these two language models have encoded two meaningfully different causal models. We believe this same procedure can underpin more thorough auditing of LLMs for misinformation or discrimination, enabling, for example, path-specific counterfactual fairness analysis [36, 10].
6 Conclusion and Future Work
In this work, we have introduced sequence-driven structural causal models (SD-SCMs)as a framework for LLM-based data generation with controllable causal structure, allowing for observational, interventional, and counterfactual data generation according to user-defined DAGs and LLM-defined structural equations. We demonstrate an important use-case for SD-SCMs by creating a benchmark for causal effect estimation. In this proof of concept, we focused on methods that assume observed confounders, but there are many other settings to explore for effect estimation, such as instrumental variables [1, 26]. Using SD-SCMs to additionally test causal discovery is of immediate interest, for example, allowing us to test whether a structure learning method can identify whether one variable is causally upstream or downstream of another [34]. Another significant area of future work is to use SD-SCMs or similar as a means of specifying causal structure over sequential data during learning [29]; rather than use an already-trained language model to generate effects, a model can be trained or fine-tuned to handle tasks that require causal reasoning, including complex confounding and sequential decision making.
In short, we believe SD-SCMs can serve as a stepping stone for any application that would benefit from sequential data with controllable causal structure, opening the door to several types of applications.
Acknowledgments
This research was supported in part by National Science Foundation (NSF) award No. 1922658 and the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI). We are grateful to Roman Mutel and Yulia Maksymiuk for involvement in early discussions of this work and especially grateful to Julia Stoyanovich for helpful feedback on data generation.
Appendix A Full description of the breast cancer SD-SCMs
The 14 covariates in the breast cancer SD-SCMs are defined generally below. For each covariate, 10 different phrasings are considered, resulting in a sample space of $10^{14}$ possible sequences. For example, for the covariate $\tilde{u}_{1}$ that represents ‘age,’ with $\Omega_{\tilde{u}_{1}}=(25,35,45,55,65,75,85)$, two possible phrasings are:
  1. 1.
    A $\tilde{u}_{1}$-year-old woman seeks consultation at the oncology clinic after being recently diagnosed with invasive breast cancer.
  2. 2.
    At the oncology clinic, a $\tilde{u}_{1}$-year-old woman is evaluated following a recent diagnosis of invasive breast carcinoma.
We consider 50 different variations of this SD-SCM, where the sample space for a given SD-SCM is defined by choosing a randomly sampled phrasing from among the possible phrasings for each of the covariates. For each of the 50 SD-SCMs, 20 datasets (each of size 1000) are sampled. Each covariate and corresponding (ordered) sample space is defined as follows.
  1. 1.
    $\tilde{u}_{1}$: age, $\Omega_{\tilde{u}_{1}}=$ (25, 35, 45, 55, 65, 75, 85)
  2. 2.
    $\tilde{u}_{2}$: medical condition, $\Omega_{\tilde{u}_{2}}=$ (hypertension, type 2 diabetes mellitus, hyperlipidemia, osteoporosis)
  3. 3.
    $\tilde{u}_{3}$: medications, $\Omega_{\tilde{u}_{3}}=$ (lisinopril, metformin, atorvastatin, calcium carbonate)
  4. 4.
    $\tilde{u}_{4}$: menopausal status, $\Omega_{\tilde{u}_{4}}=$ (pre-menopausal, post-menopausal)
  5. 5.
    $\tilde{x}_{1}$: type of carcinoma, $\Omega_{\tilde{x}_{1}}=$ (invasive ductal carcinoma (IDC), invasive lobular carcinoma, medullary carcinoma, tubular carcinoma)
  6. 6.
    $\tilde{x}_{2}$: histology grade, $\Omega_{\tilde{x}_{2}}=$ (grade 1, grade 2, grade 3)
  7. 7.
    $\tilde{x}_{3}$: genetic mutation 1, $\Omega_{\tilde{x}_{3}}=$ (TP53, PIK3CA, BRCA1, BRCA2)
  8. 8.
    $\tilde{x}_{4}$: genetic mutation 2, $\Omega_{\tilde{x}_{4}}=$ (TP53, PIK3CA, BRCA1, BRCA2)
  9. 9.
    $\tilde{x}_{5}$: level of hormone receptor expression, $\Omega_{\tilde{x}_{5}}=$ (low, moderate, high)
  10. 10.
    $\tilde{x}_{6}$: genomic instability score, $\Omega_{\tilde{x}_{6}}=$ (low, high)
  11. 11.
    $\tilde{x}_{7}$: chromosomal aberration strength, $\Omega_{\tilde{x}_{7}}=$ (significant, minor)
  12. 12.
    $\tilde{x}_{8}$: HER2 status, $\Omega_{\tilde{x}_{8}}=$ (positive, negative)
  13. 13.
    $\tilde{t}$: PD-L1 expression levels, $\Omega_{\tilde{t}}=$ (low, high)
  14. 14.
    $\tilde{y}$: therapy plan, $\Omega_{\tilde{y}}=$ (initiate an aromatase inhibitor therapy, administer a combination of a PARP inhibitor and chemotherapy, start a regimen of trastuzumab and pertuzumab, begin treatment with a checkpoint inhibitor such as pembrolizumab)
The following is an example sequence randomly sampled from one possible choice of phrasings:
A 65-year-old woman comes to the oncology department with a recent diagnosis of invasive breast carcinoma. Her prior medical history includes hyperlipidemia. This has been managed with lisinopril. This post-menopausal woman has no prior history of breast surgeries or hormone replacement therapy. Following a detailed assessment with imaging and biopsy, the results from the biopsy were analyzed and disclosed the following. The pathology report indicates the tumor is tubular carcinoma. The tumor’s histology is rated as grade 3. The tumor shows an elevated mutation burden, with particular mutations detected in the BRCA2 gene in addition to the TP53 gene. The immunohistochemistry results display robust positive staining for estrogen receptor (ER) and progesterone receptor (PR), indicating high levels of expression. The level of genomic instability in the tumor is described as low. This implies that chromosomal aberrations are minor. Immunohistochemistry reveals HER2 as negative while FISH confirms that HER2 amplification is not present. Programmed death-ligand 1 (PD-L1) expression in the tumor is low with no distant metastases found in the imaging studies. Considering the comprehensive findings and the patient’s health and treatment history, which treatment strategies are most suitable for this patient? The best option is to begin treatment with a checkpoint inhibitor such as pembrolizumab.
With each sample space indexed according to the order of their values above (with indexes starting at zero), the above text sequence corresponds to the observation
$(\tilde{u}_{1},\ldots,\tilde{u}_{4},\tilde{x}_{1},\ldots,\tilde{x}_{8},\tilde{t},\tilde{y})=(4,2,0,1,3,2,3,0,2,0,1,1,0,3).$
The full set of possible examples and code to generate this SD-SCM is available in our repository at https://github.com/lbynum/sequence-driven-scms.
Appendix B Additional estimation results
Figure 9: PEHE across all methods that estimate CATE, shown in standard deviation units of the outcome $\tilde{y}$.
Figure 10: PEHE across all methods that estimate CATE, shown in units of ITE standard deviation.
Figure 9 shows standardized Precision in Estimating Heterogeneous Effects (PEHE) [27], which is the RMSE of the CATE predictions across the different observed values of $x$, i.e.,
$\text{PEHE}_{j}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\hat{\tau}_{i}-\tau_{i})^{2}}$
for a dataset $D_{j}=\{\mathbf{u}_{i},\mathbf{x}_{i},t_{i},y_{i}\}_{i=1}^{n}$ where $\hat{\tau}_{i}$ is the estimated ITE for unit $i$ and $\tau_{i}$ is the true ITE. We standardize PEHE using the empirical standard deviation $\hat{\sigma}_{j}$ of the outcomes $\{y_{i}\}_{i=1}^{n}$ in each dataset, i.e.,
[…]
Figure 10 shows the same metric standardized instead using the (much smaller) standard deviation of the ITE.